# sample.bib

### v1

@@ Abstract
The reproduction of state-of-the-art multimodal LLM pre-training faces barriers at every stage of the pipeline, including high-quality data filtering, multimodal data mixture strategies, sequence packing techniques, and training frameworks. We introduce Open-Qwen2VL, a fully open-source 2B-parameter Multimodal Large Language Model pre-trained efficiently on 29M image-text pairs using only 442 A100-40G GPU hours. Our approach employs low-to-high dynamic image resolution and multimodal sequence packing to significantly enhance pre-training efficiency. The training dataset was carefully curated using both MLLM-based filtering techniques (e.g., MLM-Filter) and conventional CLIP-based filtering methods, substantially improving data quality and training efficiency.
@@ 1 Introduction
Ideflics Closed Closed Closed Closed Closed Open Open
@@ 2 DataComp DFN 15M huggingface:apf1/datafilteringnetworks_2b
4 DataComp MLMFilter & DFN 19.9M -
@@ 2.3 Multimodal Sequence Packing
The algorithm for the multimodal sequence packing is presented in Algorithm 1. Since we download and pack all image-text caption data into webdataset format and each webdataset tar file contains exactly 10k image-text caption data, the proposed multimodal sequence packing intends to regroup such 10k pairs into several 4k multimodal sequences.
The multimodal sequence packing involves three major steps: computing the multimodal length of each image-text caption sample, regroup data into several bins in which the total length of each bin is closest to 4096, and concatenate the input_ids vectors and pillow-format images. We adopt First-fit-decreasing (FFD) bin packing algorithm (johnson1973near) to pack each image-text caption data into several bins. We also follows LLaVA to insert an <image> placeholder token at the beginning of each image-text caption. We use the default <|im_end|> token as the separator between each image caption text.

### v2

@@ Abstract
The reproduction of state-of-the-art multimodal LLM pre-training faces barriers at every stage of the pipeline, including high-quality data filtering, multimodal data mixture strategies, sequence packing techniques, and training frameworks. We introduce Open-Qwen2VL, a fully open-source 2B-parameter Multimodal Large Language Model pre-trained efficiently on 29M image-text pairs using only 220 A100-40G GPU hours. Our approach employs low-to-high dynamic image resolution and multimodal sequence packing to significantly enhance pre-training efficiency. The training dataset was carefully curated using both MLLM-based filtering techniques (e.g., MLM-Filter) and conventional CLIP-based filtering methods, substantially improving data quality and training efficiency.
@@ 1 Introduction
Ideflics Open Open Open Open Open Open Open
@@ 2 DataComp DFN 15M huggingface:apf1/datafilteringnetworks_2b
4 DataComp MLMFilter & DFN 19.9M huggingface:weizhiwang/Open-Qwen2VL-Data
@@ 2.3 Multimodal Sequence Packing
The algorithm for the multimodal sequence packing is presented in Algorithm 1. Since we download and pack all image-text caption data into webdataset format and each webdataset tar file contains exactly 10k image-text caption data, the proposed multimodal sequence packing intends to regroup such 10k pairs into a set of multimodal sequences with 4096 context length.
The multimodal sequence packing involves three major steps: computing the multimodal length of each image-text caption sample, regroup data into several bins in which the total length of each bin is closest to 4096, and concatenate the input_ids vectors and pillow-format images. We adopt First-fit-decreasing (FFD) bin packing algorithm (johnson1973near) to pack each image-text caption data into several bins. We also follows LLaVA to insert an <image> placeholder token at the beginning of each image-text caption. We use the default end-of-sequence token, <|im_end|> token as the separator between each image caption text.